A five-card hand is chosen from a standard deck.
(a) How many different hands have exactly two hearts?

(b) What is the probability of getting one of the hands in (a) if a five-card hand is dealt at random?

Answer by Boreal(15235)

It is the number of ways to choose 2 hearts from 13 cards and 3 other cards from 39.
That is 13C2*39C3=78*(39*38*37)/6=712842
The number of hands is 52C5=2598960
The probability is 712842/2598960 = 0.274